3. You are an industry analyst who specializes in an industry where the market inverse demand is P = 200 − 4Q. The external marginal cost of producing the product is MCExternal = 6Q, and the internal cost is MCInternal = 12Q.a. What is the socially efficient level of output? b. Given these costs and market demand, how much output would a competitive industry produce? c. Given these costs and market demand, how much output would a monopolist produce? d. Discuss actions the government might take to induce firms in this industry to produce the socially efficient level of output.


1
Expert's answer
2021-10-28T16:22:46-0400

A) Socially efficient occurs at the point P = MCE + MCI

300-5Q = 8Q + 14Q

300 = 27Q

Q = 11.11 units

B) A competitive output is where P = MCI

300-5Q = 14Q

300 = 19Q

Q = 15.8 units

C) Monopolist produces at the point MR = MCI

300-10Q = 14Q

300 = 24Q

Q = 12.5 units

D) Impose a tax on production equal to marginal external cost of 8 dollars for every unit of output produced.